# Enhancing Binary Encoded Crime Linkage Analysis Using Siamese Network

### v1

@@ Introduction
Crime Linkage (CL) is integral to modern law enforcement, aiming to identify various types of serial offences (e.g., serial sexual assault, burglary, or robbery) and can also help link crimes to a known offender. The process focuses on the offender’s Modus Operandi (MO), a specific combination of behaviours that differentiate one offender from another, during the commission of the crimes to predict that the same individual is responsible for multiple incidents. Accurate CL assists in prioritizing investigative resources and strengthening public safety measures. Real-world crime databases, however, often contain complex data landscapes marked by high dimensionality, sparsity, and geographic-temporal heterogeneity (Grubin, Kelly, and Brunsdon 2001). While traditional methods have demonstrated efficiency in controlled scenarios, their effectiveness diminishes when deployed on datasets exhibiting these real-world complexities.
@@
Crime linkage identifies offence series and supports the detection and apprehension of prolific offenders (Burrell and Tonkin 2020), enabling resource sharing and preventing redundant efforts (Grubin, Kelly, and Brunsdon 2001). While physical evidence (e.g., fingerprints, DNA) can directly link crimes, behavioural indicators—victim approach, violence level, weapon use—offer alternative linkage when forensic traces are absent (Bennell and Jones 2005; Tonkin et al. 2017a). Behavioural Crime Linkage (BCL) rests on two theoretical pillars: behavioural consistency (similar offender actions across offences) and behavioural distinctiveness (unique patterns differentiating offenders) (Woodhams and Bennell 2014). Empirical support for these assumptions is extensive (Burrell, Costello, and Woodhams 2024). In practice, practitioners employ both individual behaviours and thematic clusters representing shared functions (Alison et al. 2011). Research has explored various multivariate linkage techniques, including taxonomic similarity metrics (Woodhams, Grant, and Price 2007), discriminant function analysis and multidimensional scaling for theme derivation (Winter et al. 2013), and other advanced methods (Burrell, Costello, and Woodhams 2024).
@@
Early crime linkage approaches relied on statistical models such as logistic regression and decision trees using predefined modus operandi (MO) features to capture behavioural consistencies (Bennell and Jones 2005; Melnyk et al. 2011). While effective for controlled scenarios, these models struggle with non-linear dependencies intrinsic to crime data. Recent advances leverage machine learning to process large datasets and uncover complex patterns (Bennell et al. 2014). Li and Qi (Li and Qi 2019) demonstrated enhanced serial-crime detection by combining natural language processing with dynamic time warping on crime narratives. However, challenges remain for sexual-offence linkage, where geographically dispersed, small samples often lead to under-exploited spatial and temporal features (Woodhams, Hollin, and Bull 2008; Grubin, Kelly, and Brunsdon 2001).
@@
With the advancement of artificial intelligence, ML methods have also been applied to crime linkage tasks. (Stalidis, Semertzidis, and Daras 2021) examined various deep learning architectures for crime classification and prediction, showing the robustness of learned methods outperforming traditional methods when working with 2D incident images as input. Similarly, (Utsha et al. 2024) also reviewed deep learning-based crime prediction models and highlighted the effectiveness of specific architectures in dealing with sparse crime data. Furthermore, (Butt et al. 2020) explored the application of spatio-temporal neural networks for predicting crime hotspots, seeking optimal configurations for crime data analysis. The recent study by (Burrell, Costello, and Woodhams 2024) presents a comprehensive review of the application of ML and the other approaches in CL. One notable approach in the learned domain CL is the use of Siamese Autoencoders. These networks optimize the similarity between linked crimes while maximizing the differences between unlinked ones, rather than relying on predefined similarity metrics. (Solomon et al. 2020) applied Siamese neural networks to crime data, using embeddings derived from textual features combined with spatial-temporal information to predict linkages between burglary cases. In this work, we extend the application of Siamese architectures by introducing a novel geographic-temporal integration approach designed to process high-dimensional binary-encoded features.
@@
$\mathcal{L}_{\text{contrast}}$ clusters cases by the same offender while separating dissimilar ones using a hybrid Euclidean-Manhattan distance metric (Tonkin et al. 2017b):
@@
where $y$ indicates case linkage (1=linked, 0=unlinked), $m=5$ is the margin parameter (empirically determined from values 1-10 for optimal clustering-separation balance), and $\alpha=1$ is the scaling factor(Liu, Wang, and Liu 2023; Ghojogh et al. 2020).
@@
Our data reduction approach addresses the inherent sparsity and high dimensionality of ViCLAS data through expert-informed feature consolidation of 446 binary dimensions with approximately 91% zero values, which creates challenges for pattern learning due to overwhelming inactive features (Lim, Abdullah, and Jhanjhi 2021).
@@
Expert-Driven Mapping Strategy. As shown in Table. 1, We developed five data reduction strategies through collaboration with NCA domain experts, each capturing different operational perspectives on behavioral crime linkage. This hierarchical framework groups semantically similar variables under abstract categories, identifying behaviorally similar cases despite surface-level differences (Woodhams, Grant, and Price 2007). Our mapping strategies were developed as follows: Map 1 was created by an NCA analyst with 20+ years experience, focusing on investigative relevance and operational utility. Map 2 refined Map 1 through consultation with forensic psychologists, emphasizing behavioral consistency principles. Map 3 merged Maps 1 and 2 by selecting more abstract variables where they diverged, creating a hybrid approach. Map 4 was developed by forensic psychology experts based on crime linkage literature, prioritizing behavioral distinctiveness. Map 5 represented a refined version of Map 4 with reduced abstraction to maintain behavioral specificity. For details of data reduction strategies, please refer to the Supplementary Materials.
@@ Appendix B Network Design Motivation
Siamese Networks for Sparse Crime Data. Our dataset presents additional challenges with extreme sparsity ( 9% of cases in series) and high dimensionality. Logistic regression, with its linear boundaries, fails to capture the complex behavioral interactions essential for crime linkage (Berk and Bleich 2013; Tollenaar et al. 2019). In contrast, Siamese networks excel in sparse, high-dimensional scenarios by learning latent representations that capture interaction effects between features (Pei, Tax, and van der Maaten 2016) and demonstrate superior performance over logistic regression under limited and highly imbalanced tabular data conditions (Basu et al. 2022)—precisely matching our ViCLAS dataset characteristics.

### v2

@@ Introduction
Crime Linkage (CL) is integral to modern law enforcement, aiming to identify various types of serial offences (e.g., serial sexual assault, burglary, or robbery) and can also help link crimes to a known offender. The process focuses on the offender’s Modus Operandi (MO), a specific combination of behaviours that differentiate one offender from another, during the commission of the crimes to predict that the same individual is responsible for multiple incidents. Accurate CL assists in prioritizing investigative resources and strengthening public safety measures. Real-world crime databases, however, often contain complex data landscapes marked by high dimensionality, sparsity, and geographic-temporal heterogeneity (Grubin et al. 2001). While traditional methods have demonstrated efficiency in controlled scenarios, their effectiveness diminishes when deployed on datasets exhibiting these real-world complexities.
@@
Crime linkage identifies offence series and supports the detection and apprehension of prolific offenders (Burrell and Tonkin 2020), enabling resource sharing and preventing redundant efforts (Grubin et al. 2001). While physical evidence (e.g., fingerprints, DNA) can directly link crimes, behavioural indicators—victim approach, violence level, weapon use—offer alternative linkage when forensic traces are absent (Bennell and Jones 2005; Tonkin et al. 2017b). Behavioural Crime Linkage (BCL) rests on two theoretical pillars: behavioural consistency (similar offender actions across offences) and behavioural distinctiveness (unique patterns differentiating offenders) (Woodhams and Bennell 2014). Empirical support for these assumptions is extensive (Burrell et al. 2024). In practice, practitioners employ both individual behaviours and thematic clusters representing shared functions (Alison et al. 2011). Research has explored various multivariate linkage techniques, including taxonomic similarity metrics (Woodhams et al. 2007), discriminant function analysis and multidimensional scaling for theme derivation (Winter et al. 2013), and other advanced methods (Burrell et al. 2024).
@@
Early crime linkage approaches relied on statistical models such as logistic regression and decision trees using predefined modus operandi (MO) features to capture behavioural consistencies (Bennell and Jones 2005; Melnyk et al. 2011). While effective for controlled scenarios, these models struggle with non-linear dependencies intrinsic to crime data. Recent advances leverage machine learning to process large datasets and uncover complex patterns (Bennell et al. 2014). Li and Qi (Li and Qi 2019) demonstrated enhanced serial-crime detection by combining natural language processing with dynamic time warping on crime narratives. However, challenges remain for sexual-offence linkage, where geographically dispersed, small samples often lead to under-exploited spatial and temporal features (Woodhams et al. 2008; Grubin et al. 2001).
@@
With the advancement of artificial intelligence, ML methods have also been applied to crime linkage tasks. (Stalidis et al. 2021) examined various deep learning architectures for crime classification and prediction, showing the robustness of learned methods outperforming traditional methods when working with 2D incident images as input. Similarly, (Utsha et al. 2024) also reviewed deep learning-based crime prediction models and highlighted the effectiveness of specific architectures in dealing with sparse crime data. Furthermore, (Butt et al. 2020) explored the application of spatio-temporal neural networks for predicting crime hotspots, seeking optimal configurations for crime data analysis. The recent study by (Burrell et al. 2024) presents a comprehensive review of the application of ML and the other approaches in CL. One notable approach in the learned domain CL is the use of Siamese Autoencoders. These networks optimize the similarity between linked crimes while maximizing the differences between unlinked ones, rather than relying on predefined similarity metrics. (Solomon et al. 2020) applied Siamese neural networks to crime data, using embeddings derived from textual features combined with spatial-temporal information to predict linkages between burglary cases. In this work, we extend the application of Siamese architectures by introducing a novel geographic-temporal integration approach designed to process high-dimensional binary-encoded features.
@@
$\mathcal{L}_{\text{contrast}}$ clusters cases by the same offender while separating dissimilar ones using a hybrid Euclidean-Manhattan distance metric (Tonkin et al. 2017a):
@@
where $y$ indicates case linkage (1=linked, 0=unlinked), $m=5$ is the margin parameter (empirically determined from values 1-10 for optimal clustering-separation balance), and $\alpha=1$ is the scaling factor(Liu et al. 2023; Ghojogh et al. 2020).
@@
Our data reduction approach addresses the inherent sparsity and high dimensionality of ViCLAS data through expert-informed feature consolidation of 446 binary dimensions with approximately 91% zero values, which creates challenges for pattern learning due to overwhelming inactive features (Lim et al. 2021).
@@
Expert-Driven Mapping Strategy. As shown in Table. 1, We developed five data reduction strategies through collaboration with NCA domain experts, each capturing different operational perspectives on behavioral crime linkage. This hierarchical framework groups semantically similar variables under abstract categories, identifying behaviorally similar cases despite surface-level differences (Woodhams et al. 2007). Our mapping strategies were developed as follows: Map 1 was created by an NCA analyst with 20+ years experience, focusing on investigative relevance and operational utility. Map 2 refined Map 1 through consultation with forensic psychologists, emphasizing behavioral consistency principles. Map 3 merged Maps 1 and 2 by selecting more abstract variables where they diverged, creating a hybrid approach. Map 4 was developed by forensic psychology experts based on crime linkage literature, prioritizing behavioral distinctiveness. Map 5 represented a refined version of Map 4 with reduced abstraction to maintain behavioral specificity. For details of data reduction strategies, please refer to the Supplementary Materials.
@@ Conclusion
Acknowledgments
This research was partially supported by funding from the National Crime Agency, UK. The authors would also like to thank the analysts from the SCAS unit at the National Crime Agency for their valuable assistance with data preparation and for providing insightful feedback on the research findings.
@@ Appendix B Network Design Motivation
Siamese Networks for Sparse Crime Data. Our dataset presents additional challenges with extreme sparsity ( 9% of cases in series) and high dimensionality. Logistic regression, with its linear boundaries, fails to capture the complex behavioral interactions essential for crime linkage (Berk and Bleich 2013; Tollenaar and others 2019). In contrast, Siamese networks excel in sparse, high-dimensional scenarios by learning latent representations that capture interaction effects between features (Pei et al. 2016) and demonstrate superior performance over logistic regression under limited and highly imbalanced tabular data conditions (Basu et al. 2022)—precisely matching our ViCLAS dataset characteristics.